suppose a confidence interval is 9.65<mean<11.35. Find the best point estimate of the population mean, find the error estimate E.
Answer #1

Solution :

Given that,

Lower confidence interval = 9.65

Upper confidence interval = 11.35

​  = (Lower confidence interval + Upper confidence interval ) / 2

= (9.65 + 11.35) / 2

= 21 / 2 = 10.5

= 10.5

Point estimate = 10.5

Margin of error = E = Upper confidence interval - ​  = 11.35 - 10.5 = 0.85

error estimate E = 0.85